Lava Lake Inc. bottles and distributes spring water. On February 11 of the current year, Lava Lake reacquired 6,000 shares of its common stock at $74 per share. On April 30, Lava Lake Inc. sold 4,200 of the reacquired shares at $80 per share. On August 22, Lava Lake Inc. sold 1,800 shares at $70 per share.

a. Journalize the transactions of February 11, April 30, and August 22. For a compound transaction, if an amount box does not require an entry, leave it blank.

Feb. 11
Apr. 30
Aug. 22

b. What is the balance in Paid-In Capital from Sale of Treasury Stock on December 31 of the current year?
$  

c. For what reasons might Lava Lake have purchased the treasury stock?
Answer #1
a
Feb 11 Treasury Stock 444000 =6000*74
      Cash 444000
Apr 30 Cash 336000 =4200*80
     Treasury Stock 310800 =4200*74
     Paid in Capital from sale of Treasury Stock 25200
Aug 22 Cash 126000 =1800*70
Paid in Capital from sale of Treasury Stock 7200
     Treasury Stock 133200 =1800*74
b
Paid-In Capital from Sale of Treasury Stock 18000 Credit =25200-7200
c
Reasons for purchase of Treasury Stock:
For reissuance to employees as a bonus according to stock purchase agreements
To support the market price of the stock
To provide shares for resale to employees